Hi,
I keep trying to figure out where I am going wrong in this equation but I cannot. The equation is...
a=b
a^2=ab
a^2-b^2=ab-b^2
(a+b)(a-b)=b(a-b)
a+b=b
a+a=a
2a=a
2=1

Answer by Edwin McCravy(20055)

a=b
 a^2=ab
 a^2-b^2=ab-b^2
 (a+b)(a-b)=b(a-b)
 a+b=b   <-- this step is wrong because you can only divide both
             sides by a number other than zero, however the first
             step is a=b which means that a-b = 0, so this step
             is incorrect because it divides the preceding step
             through by zero.   

Edwin